OpenStudy (anonymous):

Simplify: Radical 16r ^7 s^5+ 7r^2 Radical r^3s^5-7rs radical r^5s^3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sqrt{16r^7s^5}+7r^2\sqrt{r^3 s^5}-7rs \sqrt{r^5s^3}\] you mean this?

OpenStudy (anonymous):

im guessing each are different?

OpenStudy (anonymous):

need clarification

OpenStudy (anonymous):

yes i meant that \. I didn't know how to to type it

OpenStudy (anonymous):

o.k, so first look at the radical sign

OpenStudy (anonymous):

you want to grab pairs

OpenStudy (anonymous):

so you can rewrite it as \[\sqrt{16r^6rs^4s}\]

OpenStudy (anonymous):

for the first term

OpenStudy (anonymous):

so just take each individual root

OpenStudy (anonymous):

\[\sqrt{16}=4\]

OpenStudy (anonymous):

one way to look at the r^6 is to divide the exponent by 2

OpenStudy (anonymous):

you will do the same to the s^4

OpenStudy (anonymous):

Ive done all of this correct, but i keep getting the wrong answer

OpenStudy (anonymous):

we need pairs so we will have r and s inside the radical.

OpenStudy (anonymous):

o.k, what are u getting?

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

what is you final answer?

OpenStudy (anonymous):

im getting \[4r^3s^2\sqrt{rs}\]